Grade 12Physical Chemistry

If f(x) = x|x| then inverse of f(x) equals –
(where sgn(x) denotes signum function of x)
Ans: sgn(x)* [x]0.5

Harshit Bhansali
6 Years agoGrade 12

1 Answer

Arun
ApprovedApproved Tutor Answer6 Years ago
Dear Student
 
y = x^2
Now x = sqrt(y)
 
Hence inverse  = f-1 (x) = srqt (x)
 
But to cinscinsider the positive or negative sign,
We have to usw signum function.
 
Hence f-1(x) = sgn (x) srqt(x)